OpenStudy (anonymous):

Calvin has $8.80 in pennies and nickels. If there are twice as many nickels as pennies, how many pennies and nickels does Calvin have?

OpenStudy (itiaax):

Let x be the no. of pennies and let y be the no. of nickles

OpenStudy (itiaax):

Set up equations using these variables: Equation 1: y=2x Equation 2: 0.01x + 0.5y = 8.80 Now, you solve: Substiture y=2x in equation 2: = 0.01x + 0.5(2x) = 8.80 = 0.01x + 0.10x = 8.80 = 0.11x=8.80 x = 8.80/0.11 = 80 Therefore, x=80. Substiture x=8 in equation 1: y=2(80) = 160 Therefore y=160. Hence, there are 80 pennies and 160 nickels.